OpenStudy (anonymous):

Jessi rents a moving van for $75 and must pay $2 per mile. The following week, Alex rents the same van, is charged $80 for the rental and $1.50 per mile. If they each paid the same amount and drove the same number of miles, how far did they each travel? I already did this and got 310miles. Is that right? Ill show you how i set it up below.

OpenStudy (anonymous):

m for miles, 75 + 2m = 80 + 1.5m, 155 = .5m, traveled 310 miles Seem right anyone?!